# Tree-of-Code: A Tree-Structured Exploring Framework for End-to-End Code Generation and Execution in Complex Task Handling

### v1

@@ Abstract
Solving complex reasoning tasks is a key real-world application of agents. Thanks to the pretraining of Large Language Models (LLMs) on code data, recent approaches like CodeAct successfully use code as LLM agents’ action, achieving good results. However, CodeAct greedily generates the next action’s code block by relying on fragmented thoughts, resulting in inconsistency and instability. Moreover, CodeAct lacks action-related ground-truth (GT), making its supervision signals and termination conditions questionable in multi-turn interactions. To address these issues, we first introduce a simple yet effective end-to-end code generation paradigm, CodeProgram, which leverages code’s systematic logic to align with global reasoning and enable cohesive problem-solving. Then, we propose Tree-of-Code (ToC), which self-grows CodeProgram nodes based on the executable nature of the code and enables self-supervision in a GT-free scenario. Experimental results on two datasets using ten popular zero-shot LLMs show ToC remarkably boosts accuracy by nearly 20% over CodeAct with less than 1/4 turns. Several LLMs even perform better on one-turn CodeProgram than on multi-turn CodeAct. To further investigate the trade-off between efficacy and efficiency, we test different ToC tree sizes and exploration mechanisms. We also highlight the potential of ToC’s end-to-end data generation for supervised and reinforced fine-tuning.
Tree-of-Code: A Tree-Structured Exploring Framework for End-to-End Code Generation and Execution in Complex Task Handling
Ziyi Ni1,2, Yifan Li4, Ning Yang1, Dou Shen3, Pin Lv1, Daxiang Dong3, 1The Key Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences, 2School of Artificial Intelligence, University of Chinese Academy of Science 3Baidu, Inc., 4Global Innovation Exchange Institution, Tsinghua University, Correspondence: dongdaxiang@baidu.com, pin.lv@ia.ac.cn,
@@ 1 Introduction
The application of code generation techniques to complex task planning and execution has garnered significant attention Holt et al. (2024); Wen et al. (2024b); Xu et al. (2024), particularly with the emergence of CodeAct Wang et al. (2024) approaches. CodeAct moves the interaction unit from ReAct’s individual tool calls to generating code blocks with local reasoning, while uniquely using code logic and libraries. Rather than JSON Qin et al. (2023) or text Park et al. (2023), it treats code as action, utilizing LLM’s pre-trained coding skills for efficient handling of complex tasks.
However, each turn of CodeAct is based on individual actions rather than the entire program. It can not autonomously reason and generate complete code in one turn. Instead, it follows a step-by-step generation and interaction process. This is akin to the brain’s control of motor actions, processing tasks iteratively and incrementally, with the basal ganglia supporting step-granularity execution and the cerebellum refining and sequencing for the complete complex motor task Baladron et al. (2023).
On the one hand, the fragmented and stalled thinking in code can hinder a thorough understanding of the logical chains embedded within it, potentially leading to redundant code Wang et al. (2023); Guo et al. (2024). Repeatedly integrating all prior thoughts causes context overload due to long histories, making it more prone to accumulating significant model hallucinations Ji et al. (2023).
On the other hand, solving complex problems can have multiple solutions Mialon et al. (2023), such as calling tools in different sequences. LLMs also tend to explore randomly, resulting in varied solutions, making it difficult to set a standard answer for each step. Long action sequences over multiple turns only provide sparse rewards Xu et al. (2023). When using these trajectories as training data for supervised fine-tuning (SFT), they can only be combined into an overall program response as one sample Wang et al. (2024). Applying reinforcement learning is challenging due to the lack of necessary process supervision Zelikman et al. (2024), leading to fundamental issues.
Therefore, we are considering how to use task-level feedback as a supervision signal for each turn. Additionally, how can we design a framework that incorporates reflection and refinement based on environmental feedback to create real multi-turn trajectory data?
Here, we are inspired by the Dorsolateral Prefrontal Cortex (DPC), which is involved in high-level global cognitive circuits Kaller et al. (2015), to simulate how programmers form systematic representations and utilizations of code. We first propose an end-to-end code reasoning and generation paradigm, dubbed CodeProgram. In this way, the final answer can serve as a direct evaluation metric.
To utilize environmental feedback from code execution, we further develop a framework called Tree-of-Code (ToC). In this method, task-level CodePrograms serve as nodes, forming a self-expanding exploration tree driven by the verifiability of code execution. This differs from "Tree-of-Thoughts" (ToT) Yao et al. (2024a), which enhances "Chain-of-Thought" (CoT) Wei et al. (2022) by exploring varied approaches within the same solution. In contrast, we generate diverse solutions through random settings, with each node representing a complete solution. This concept is akin to a Code "Random Forest" Rigatti (2017). Experimental results on two types of multi-turn, multi-tool, complex task datasets, using ten LLMs, demonstrate that ToC outperforms CodeAct in both accuracy and efficiency. The core contributions of this paper are summarized as follows:
@@
    We introduce CodeProgram, an end-to-end paradigm designed to continuously generate complete code solutions using necessary tools. The purpose and its advantages are analyzed.
@@
    We present Tree-of-Code (ToC), which self-grows CodeProgram nodes by leveraging the executable nature of the code. By exploring random settings, ToC enhances performance in solving complex tasks through ensembling.
@@
    Empirical validation on two complex task datasets, M3ToolEval and API-Bank, demonstrates significant improvements in both problem-solving performance and efficiency.
@@ 2 Design Motivation
In industry, complex tasks requiring multiple tools and function calls, are typically driven by open-ended user queries. This creates two key challenges: (1) For zero-shot queries, it is challenging to pre-obtain task-level ground-truth (GT). Manual annotation or assigning different rewards to responses is required for subsequent SFT Chung et al. (2024) or reinforced fine-tuning (ReFT) Luong et al. (2024). Moreover, without GT, the termination criteria become unclear. (2) Multi-turn interactions lack a fixed trajectory, making it difficult to define the process supervised signals Luo et al. (2024). Current methods often rely on a ’judge’ model to evaluate whether the user’s needs from the task query are met at each step Chen et al. (2024); Li et al. (2024a). However, each evaluation demands strong analytical and reasoning skills from the model, making it costly and time-consuming. Existing methods deliberately avoid these challenges by assuming GT is known, matching task-level GT with action-related outcomes at each step. The process stops if they match, or continues until the step limit is reached. The tool agent in Figure 1 illustrates this.
We aim to explore whether it is possible to develop a method that can self-provide supervision and termination criteria at each step while approximating GT in a GT-free scenario.
Unlike cerebellum-controlled step-by-step motor tasks, we learn from the DPC to treat each turn as a complete task. By iteratively approaching the feasible region, we collect a batch of feasible solutions and then consider the optimal one. Inspired by it, we represent the program agent in Figure 1 to simulate the cognitive process of senior programmers. When working on a project, they start with global planning, complete the entire code, and iteratively debug until no errors are reported, leveraging the systematic logic and environmental feedback inherent to the code project.
Figure 2: Illustration of CodeProgram.
3 CodeProgram
A block of code in CodeAct’s actions corresponds to the eyes or mouths of the smiley face in Figure 1, when faces represent complex tasks. We propose CodeProgram, which can draw the complete face in one turn. Figure 2 illustrates how it works. Specifically, code serves as a bridge, aligning with natural language reasoning and connecting it to execution outcomes in the environment. By decoupling the reasoning process from code execution, we achieve flexibility while ensuring consistency.
3.1 Code as Reasoning
Code generation plays a crucial role in the concept of "code-as-reasoning," where the process of writing code itself reflects a reasoning process.
Global reasoning is required to guide complete code generation in a single end-to-end process. This enables the seamless integration of various reasoning methods for large language models (LLMs), such as prompt engineering Chen et al. (2023a), Chain-of-Thoughts (CoT) Wei et al. (2022), Tree-of Thoughts (ToT) Yao et al. (2024a), in-context learning Kojima et al. (2022), self-reflection  Zhang et al. (2024), and System2 reasoning Frankish (2010); OpenAI (2024b); Yao et al. (2024b). Furthermore, longer chains of thought have consistently been shown to improve task performance Wei et al. (2022); Zelikman et al. (2024).
@@ 3.1 Code as Reasoning
Figure 3: An Overview of CodeAct and ToC. (a) CodeAct regards code as action with step-by-step reasoning. (b) ToC applies execution-based reflection in the tree structure, where each node generates end-to-end code with global planning as its thoughts. At each layer, different nodes are executed in parallel; if executed successfully, they are collected for voting. Yellow boxes with dashed borders indicate invalid nodes that fail execution. Both red and green boxes represent valid nodes: red boxes are discarded through LLM voting, while green boxes are finally accepted. The query is "Find nearby restaurants within 1km of San Francisco" from API-Bank level-3 dataset.
3.2 Two Helper Tools
CodeProgram struggles when LLMs must rely on tool outputs to determine the next steps. For example, we can only provide the final summary answer based on all tool outputs; in web browsing tasks, the next action is determined only after the page content is viewed. To maintain end-to-end flow, we introduce two additional functions that call the LLM into our code: a general res_handler, which defines a prompt to generate results that meet the prompt requirements for final summarization, and a specific next_action for web tasks, which decides the next action from a given set of possible browsing actions based on the page content, visited URLs, and task query. Their tool descriptions and functions are shown in Appendix B. They help better understand the semantic relationships between tools, ensuring a smooth and cohesive sequence of tool calls during code generation.
3.3 Execution Outcome can be Label
The code solution is task-level, and its execution outcome is a self-provided annotation that can be directly used as labels. Specifically, it includes supervision signals to select the "successfully executed" samples for SFT, and various rich comments (such as specific results or error messages) that can be quantified as rewards for ReFT, as long as we repeat the CodeProgram in different settings multiple times to receive feedback. In this way, the code acts as a verifier. This idea inspires us to build a single-layer multi-node Tree-of-Code (ToC). Thanks to the task-level granularity, the code’s execution outcomes align with the task query and the thought-code output, facilitating the generation of valuable training data, as the thoughts, code, and execution-based labels are tightly synchronized.
In summary, CodeProgram is an annotation-free, end-to-end generation approach well-suited for producing large-scale training data, increasing efficiency, and improving task-solving performance.
@@ 4 Tree-of-Code Method
Following the design motivation in Figure 1, we need to collect all successfully executed solutions and identify the one closest to the GT. CodeProgram introduced in Section 3 has allowed us to achieve two key goals: (1) to directly reflect on and refine the task-level code, and (2) to use its execution results as terminal criteria. We now propose an execution-based, self-growing, and self-filtering tree, with CodeProgram as its nodes.
@@ 4.1 Overview of Tree-of-Code
We represent the ToC framework as $\textbf{{T}}=(\textbf{{N}, {S}})$, where N denotes a set of nodes ($N$), and S represents the stems (unidirectional arrows in Figure 3) , modeling the reflection reasoning process of LLMs when expanding the nodes. The overview of ToC and how it works is illustrated in Figure 3. Let $L$ denote the max depth, $l$ the layer index, $M$ the expanded layer’s max width, $m$ the node index, $l\in\{1,\dots,L\}$, $m\in\{1,\dots,M\}$. We use $T$ for the thoughts of the $N$, $C$ for code, and $E$ for its execution result. The next-layer $N$ is denoted as:
@@ 4.2 Tree Expansion
We initialize from a root node and recursively expand the tree. The expansion process follows: (1) The breadth-first search (BFS) strategy is applied, with each parent node branching into $M$ child nodes. (2) Whether the node continues to grow depends solely on the evaluation of its own execution state (success or failure). For each $N_{l}$,
$\left\{\begin{array}[]{ll}\text{{stop and collect},}&\text{{if} }E_{l}\neq\text{{None or error}},\\ \text{{grow} }\textbf{{N}}_{(l+1)},&\text{{otherwise}.}\end{array}\vspace{-3pt}\right.$
@@
Execution-based Reflection. We can not guarantee that the end-to-end code solution will be correct on the first attempt. Treating task-level execution errors as continuation signals, we propose execution-based reflection, which enables LLMs to self-reflect, identify errors, refine thoughts, and improve code through prompting, significantly enhancing problem-solving. The prompt for reflection is shown in Appendix A.2.1.
As long as execution fails, self-reflection continues iteratively, generating new nodes. This allows the branch to grow as a data sample, with each node in the trajectory providing supervision signals. Since these supervision signals are inherently embedded within the CodeProgram node, the growth process is self-driven. Therefore, the whole tree is end-to-end generated. Figure 4 illustrates an example of a branch of ToC.
Additionally, our flexible tree-structured framework allows for the integration of any reflection method in end-to-end code generation.
@@
Research Renze and Guven (2024) has shown that performance benefits from diverse perspectives of error identification, which encourages models to generate multiple solutions.
To enhance the diversity of ToC, we introduce randomness into the expansion process by varying LLMs and prompts, inspired by the random forest Rigatti (2017). At the system level, different LLMs are randomly explored from our LLM list, which will be introduced in Section 5.1, with a consistent temperature setting of 0.1. At the instruction level, prompts are randomly selected from a diverse pool, designed through self-evolution and human crafting phases. In the first phase, we used our ten LLMs to create ten various prompts based on the prompt evolution with the root prompt (see Appendix A). The prompt for prompt evolution requires maintaining consistent core content with the root prompt while encouraging orthogonal or divergent expressions. Then, we manually selected six distinct prompts, randomly applying one or more of the following modifications: (1) Adding more detailed usage examples (beyond just printing "Hello world") to three prompts; (2) Adjusting their format by adding line breaks and indentation; (3) Randomly rearranging components, such as the reflection part, usage examples, role instructions, tool descriptions, and the chat history.

### v2

# Tree-of-Code: A Self-Growing Tree Framework for End-to-End Code Generation and Execution in Complex Tasks

@@ Abstract
Solving complex reasoning tasks is a key real-world application of agents. Thanks to the pretraining of Large Language Models (LLMs) on code data, recent approaches like CodeAct successfully use code as LLM agents’ action, achieving good results. However, CodeAct greedily generates the next action’s code block by relying on fragmented thoughts, resulting in inconsistency and accumulative hallucination. Moreover, CodeAct lacks action-related ground-truth (GT), making its supervision signals and termination conditions questionable in multi-turn interactions. To address these issues, we propose Tree-of-Code (ToC), a self-growing framework that generates nodes through self-supervision, incorporating prompt and model exploration in a GT-free setting. Each node employs CodeProgram, an end-to-end code generation paradigm that aligns executable code logic with global reasoning. This approach uses task-level execution success as both node validity and stop-growing flags, bypassing process supervision to enable online applications. Experiments on two datasets with ten popular zero-shot LLMs show that ToC boosts accuracy by nearly 20% over CodeAct with fewer than 1/4 turns. To further investigate the trade-off between efficacy and efficiency, ablation studies on different ToC tree sizes and exploration mechanisms validate ToC’s superiority.
Tree-of-Code: A Self-Growing Tree Framework for End-to-End Code Generation and Execution in Complex Tasks
Ziyi Ni1,2,thanks: These authors contributed equally to this work.thanks: This work was conducted at Qianfan AppBuilder Group during the Baidu AI Cloud Summer Camp internship., Yifan Li4,11footnotemark: 122footnotemark: 2, Ning Yang1, Dou Shen3, Pin Lv1,‡, Daxiang Dong3,‡, 1The Key Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Science 3Baidu, Inc. 4Global Innovation Exchange Institution, Tsinghua University ‡ Correspondence: dongdaxiang@baidu.com, pin.lv@ia.ac.cn
@@ 1 Introduction
The application of code generation techniques to complex task planning and execution has garnered significant attention Holt et al. (2024); Wen et al. (2024a); Xu et al. (2024b), particularly with the emergence of CodeAct Wang et al. (2024) approaches. CodeAct moves the interaction unit from ReAct’s individual tool calls to generating code blocks with local reasoning while leveraging code logic and libraries. Rather than JSON Qin et al. (2023) or text Park et al. (2023), it treats code as action, utilizing LLM’s pre-trained coding skills for efficient handling of complex tasks.
However, CodeAct treats each turn as an individual action rather than addressing the entire program, following a step-by-step generation process. While this approach may seem explorative, it has four critical limitations: (I). CodeAct assumes that the ground truth (GT) is known and uses GT matching as a termination criterion, which is unrealistic and unfeasible. (II). Fragmented thinking is inefficient. For simple problems, stalled thinking is not only unnecessary but also disrupts the logical chains in the code Wang et al. (2023); Guo et al. (2024). Moreover, as the number of turns increases, repeatedly integrating prior thoughts causes context overload, heightening model hallucinations Ji et al. (2023), increasing computational cost. (III). CodeAct lacks exploration of diverse reasoning paths. While it supports multi-turn interactions, it follows a single reasoning process. In contrast, solving complex problems often has multiple solutions Mialon et al. (2023), where different approaches can branch from different points, making it difficult to set a standard answer for each turn. (IV). Generated trajectory data is hard to reuse. When using these trajectories for supervised fine-tuning (SFT), they cannot be directly combined into a single program response Wang et al. (2024). Reinforcement learning is also challenging due to the lack of process supervision Zelikman et al. (2024), leading to fundamental issues.
Since defining and obtaining supervision signals for intermediate states is challenging, we propose using task-level feedback directly, treating task completion as a single step. We introduce CodeProgram, an end-to-end code reasoning and generation paradigm, as a ‘turn,’ where the only environmental supervision is execution success. To incorporate reflection and exploration, we design an outcome-driven refinement framework, Tree-of-Code (ToC), that enables multi-turn interactions with diverse solutions exploring the model and prompt pools as tree branches, where task-level CodePrograms serve as the nodes. The final output is determined by voting on the collected nodes, selected based on their successful execution. It’s important to note that, in this paper, a ‘turn’ refers to a single action of code generation. For our CodeProgram, a turn involves completing the entire program, rather than just a single task step (as in CodeAct).
Although ToC’s name and structure are similar to "Tree-of-Thoughts" (ToT) Yao et al. (2024), their meanings fundamentally differ. Our concept might be closer to a Code "Random Forest" Rigatti (2017). While ToT enhances "Chain-of-Thought" (CoT) Wei et al. (2022) by exploring different thoughts within the same solution, ToC explores multiple distinct program solutions. In other words, each node in ToC represents a complete solution, and the tree as a whole captures different iterative optimizations (depth) across a variety of complete solutions (breadth). The core contributions of this paper are summarized as follows:
@@
    We propose a self-growing Tree-of-Code (ToC) structure that automatically reflects and explores diverse, complete solution nodes without labeled data, facilitating complex tasks in multi-tool online scenarios.
@@
    Each node in ToC, called a CodeProgram, is generated end-to-end. We are the first to define process-level supervision at the task-outcome level using execution success.
@@
    Extensive experiments and ablation studies on two multi-tool, complex task datasets with ten models, demonstrate that ToC significantly enhances problem-solving accuracy and efficiency in real-world, zero-shot scenarios.
@@ 2 Design Motivation
Figure 2: An Overview of CodeAct and ToC. (a) CodeAct regards code as action with step-by-step reasoning. (b) ToC applies execution-based reflection in the tree structure, where each node (CodeProgram) generates end-to-end code with global planning as its thoughts. At each layer, nodes are executed in parallel; if executed successfully, they are collected for voting. Note that the process supervision relies solely on the node’s execution success or failure, rather than on the specific content executed (whether correct or incorrect), which would require pre-known labels. The query is "Find nearby restaurants within 1km of San Francisco" from API-Bank level-3 dataset.
In industry, complex tasks requiring multiple tools and function calls, are typically driven by open-ended user queries. This creates two key challenges: (1) For zero-shot queries, it is unrealistic to pre-obtain task-level ground-truth (GT), which is required for SFT Chung et al. (2024) or reinforced fine-tuning (ReFT) Luong et al. (2024). Moreover, without GT, the termination criteria become unclear. (2) Multi-turn interactions lack a standard trajectory, making it difficult to define the process supervised signals Luo et al. (2024). Current methods often rely on ’LLM-as-judge’ to evaluate whether the user’s needs are met at each step Chen et al. (2024); Li et al. (2024a). However, it would require an API call after every step to check progress, ultimately increasing both time and token costs. Besides, the evaluation without objective signals demands strong analytical and reasoning skills from LLMs. Existing methods deliberately avoid these challenges by assuming GT is known Wang et al. (2024), matching task-level GT with action-related outcomes at each step, like the tool agent in Figure 1: the interaction turn stops only if they match, or continues until the step limit is reached.
Since intermediate states are absent, if possible, why not treat each complete end-to-end execution as an atomic state? By iteratively exploring feasible solutions through parallel executions, we first collect a batch of solutions, and then determine the optimal one, as shown by the program agent in Figure 1. This idea inspires our node outcome-driven reflection system specifically designed for multi-tool interaction in real-world environments. Our key contribution is a self-growing framework enabling LLM agents to autonomously interact with code through zero-shot learning without GT supervision, whose implementation details will be subsequently presented.
@@ 3 Tree-of-Code Method
Following the design motivation, we need to collect all valid solutions and identify the one closest to the GT. By treating each tree node as a complete task-level solution and exploring different nodes for breadth while deepening through iterative refinement, we propose ToC (Tree-of-Code), an execution-based, self-growing, and self-filtering tree for handling real-world complex tasks.
@@ 3.1 Overview of Tree-of-Code
We represent the ToC framework as $\textbf{{T}}=(\textbf{{N}, {S}})$, where N denotes a set of nodes ($N$), and S represents the stems (unidirectional arrows in Figure 2) , modeling the reflection reasoning process of LLMs when expanding the nodes. The overview of ToC and how it works is illustrated in Figure 2. Let $L$ denote the max depth, $l$ the layer index, $M$ the expanded layer’s max-width, $m$ the node index, $l\in\{1,\dots,L\}$, $m\in\{1,\dots,M\}$. We use $T$ for the thoughts of the $N$, $C$ for code, and $E$ for its execution result. The next-layer $N$ is denoted as:
@@
3.2 Tree Node Generation
Unlike tool agents like CodeAct, which treat each intermediate action and environmental feedback as a step, each node in our ToC represents a complete task, effectively increasing the granularity of task handling at each layer.
In other words, a single tree node (one turn) is equivalent to multiple turns of CodeAct, with both being directly comparable and serving the same purpose (final response), significantly improving efficiency. We refer to this end-to-end code reasoning and generation paradigm as CodeProgram. Figure 4 illustrates how it works.
Specifically, the end-to-end code in CodeProgram serves as a bridge, aligning with natural language reasoning and execution outcomes in the environment. Besides, by decoupling the reasoning process from code execution, we achieve flexibility while ensuring consistency.
3.2.1 Code as Reasoning
On the one hand, CodeProgram leverages the concept of "code-as-reasoning" to generate code, where the process of writing code itself mirrors the reasoning process.
On the other hand, global reasoning is essential for guiding CodeProgram’s complete code generation in a single end-to-end flow. This approach enables the seamless integration of various reasoning techniques for large language models (LLMs), such as prompt engineering Chen et al. (2023a), Chain-of-Thoughts (CoT) Wei et al. (2022), Tree-of-Thoughts (ToT) Yao et al. (2024), in-context learning Kojima et al. (2022), self-reflection Zhang et al. (2024), and System2 reasoning Frankish (2010); OpenAI (2024b). Additionally, longer chains of thought have consistently been shown to enhance task performance Zelikman et al. (2024).
Figure 4: Illustration of the CodeProgram in ToC.
3.2.2 Two Helper Tools
CodeProgram struggles with environmental exploration when LLMs must rely on tool outputs to determine the next steps. For instance, in web browsing tasks, the next action can only be decided after viewing the page content, and a final summary answer can only be provided after considering all tool outputs. Thus, to maintain end-to-end flow, we introduce two functions: a general res_handler, which defines a prompt to generate results that meet the prompt requirements for final summarization, and a specific next_action for web tasks, which decides the next action from a given set of possible browsing actions based on the page content, visited URLs, and task query. Their tool descriptions and functions are shown in Appendix B.
They help better understand the semantic relationships between tools, ensuring a smooth, cohesive sequence of tool calls during code generation. In the Appendix B.3, we also provide an example demonstrating how these helper tools work.
3.2.3 Execution Outcome as Process Label
The code solution is task-level, and its execution outcome is a self-provided annotation that can be directly used as labels. Note that we focus solely on task execution success, using a simple true/false label to filter feasible solutions and approximate more effective ones. This label is weak but available, simple, and useful—unlike pre-known GT or correctness judgments.
Benefiting from our end-to-end paradigm (a direct, complete task-level response to a single query), we can select "successfully executed" samples for SFT and use various rich comments (such as specific results or error messages) as rewards for ReFT by repeating the CodeProgram in different settings (i.e., multi-nodes). In this context, the code acts as a verifier. This verification-then-refinement concept also inspires the development of a multi-layer Tree-of-Code (ToC).
Thanks to task-level granularity, the code’s execution outcomes align with both the task query and the thought-code output, enabling the generation of valuable data for potential future training.
@@ 3.3 Tree Expansion
We initialize from a root node and recursively expand the tree. The expansion process follows: (1) The breadth-first search (BFS) strategy is applied, with each parent node branching into $M=3$ child nodes. (2) Whether the node continues to grow depends solely on the evaluation of its own execution state (success or failure). For each $N_{l}$,
$\left\{\begin{array}[]{ll}\text{{stop and collect},}&\text{{if} }E_{l}\neq\text{{None or error}},\\ \text{{grow} }\textbf{{N}}_{(l+1)},&\text{{otherwise}.}\end{array}\right.$
@@
Execution-based Reflection. We can not guarantee that one node solution will be correct on the first attempt. Treating task-level execution errors as continuation signals, we propose execution-based reflection, which enables LLMs to self-reflect, identify errors, refine thoughts, and improve code. As long as execution fails, self-reflection continues iteratively, generating next-layer new nodes. The prompt for reflection is shown in Appendix A.2.1.
This also allows the branch to grow into deeper layers, where each node in the trajectory provides process supervision signals based on its outcome. Note that the definition of ’turn’ is equivalent to that of ’layer’; both terms carry the same meaning. Since these supervision signals are inherently embedded within the CodeProgram node, the growth process is self-driven. Therefore, the whole tree is end-to-end generated.
Figure 3 shows an example of a branch of ToC while Figure 5 demonstrates execution-based reflection and tree expansion. Additionally, our flexible tree-structured framework allows for the integration of any reflection method for tree expansion.
@@
Research Renze and Guven (2024) has shown that performance benefits from diverse perspectives of error identification, which encourages models to generate multiple solutions (ie. nodes in ToC).
To enhance the diversity of ToC, we introduce randomness into the expansion process by varying LLMs and prompts, inspired by the random forest Rigatti (2017). At the system level, different LLMs from our list, introduced in Section 1, are explored randomly with a consistent temperature setting of 0.1. At the instruction level, prompts are randomly selected from a diverse pool, created through self-evolution and human crafting.
The random exploration mechanisms operate at each node individually, while the prompt pool is created just once for the entire system.
Specifically, we used ten LLMs to generate ten diverse prompts through prompt evolution from the root prompt (see Appendix A). The evolution process ensures the core content remains consistent while promoting orthogonal or divergent expressions. Six distinct prompts were manually selected and the following modifications were then applied: (1) adding detailed usage examples (beyond just printing "Hello world") to three prompts; (2) adjusting the format with line breaks and indentation; (3) randomly rearranging components, including the reflection part, usage examples, role instructions, tool descriptions, and chat history.